​Dobosh Corporation has provided the following information:

Dobosh Corporation has provided the following information:




Cost per

Unit

Cost per Period


Direct materials

$7.05



Direct labor

$3.65



Variable manufacturing overhead

$1.60



Fixed manufacturing overhead


$113,400


Sales commissions

$1.50



Variable administrative expense

$0.55



Fixed selling and administrative expense


$36,450

Required:

a. For financial reporting purposes, what is the total amount of product costs incurred to make 9,000 units?

b. For financial reporting purposes, what is the total amount of period costs incurred to sell 9,000 units?

c. If 10,000 units are sold, what is the variable cost per unit sold?

d. If 10,000 units are sold, what is the total amount of variable costs related to the units sold?

e. If 10,000 units are produced, what is the total amount of manufacturing overhead cost incurred?

f. If the selling price is $21.60 per unit, what is the contribution margin per unit sold?

g. If 8,000 units are produced, what is the total amount of direct manufacturing cost incurred?

h. If 8,000 units are produced, what is the total amount of indirect manufacturing costs incurred?

i. What incremental manufacturing cost will the company incur if it increases production from 9,000 to 9,001 units?

16) Saxbury Corporation's relevant range of activity is 3,000 units to 7,000 units. When it produces and sells 5,000 units, its average costs per unit are as follows:




Average Cost per Unit


Direct materials

$5.30


Direct labor

$3.65


Variable manufacturing overhead

$1.50


Fixed manufacturing overhead

$3.90


Fixed selling expense

$0.75


Fixed administrative expense

$0.60


Sales commissions

$0.50


Variable administrative expense

$0.50


Required:

a. For financial reporting purposes, what is the total amount of product costs incurred to make 5,000 units?

b. For financial reporting purposes, what is the total amount of period costs incurred to sell 5,000 units?

c. If 6,000 units are sold, what is the variable cost per unit sold?

d. If 6,000 units are sold, what is the total amount of variable costs related to the units sold?

e. If 6,000 units are produced, what is the average fixed manufacturing cost per unit produced?

f. If 6,000 units are produced, what is the total amount of fixed manufacturing cost incurred?

g. If 6,000 units are produced, what is the total amount of manufacturing overhead cost incurred? What is this total amount expressed on a per unit basis?

h. If the selling price is $22.90 per unit, what is the contribution margin per unit sold?

i. If 4,000 units are produced, what is the total amount of direct manufacturing cost incurred?

j. If 4,000 units are produced, what is the total amount of indirect manufacturing cost incurred?

k. What incremental manufacturing cost will the company incur if it increases production from 5,000 to 5,001 units?
Answer #1

Product costs are the cost which are incurred on making of the product

A.total product costs incurred in making 9,000 units = direct material cost + direct labor cost + varible overhead cost

=(7.05+3.65+1.60)*9000

= $110,700

B. Period costs = fixed manufacturing overhead + sales commission + variable administrative expenses+ fixed administrative expenses

= 113,400 + (1.5+0.55)*9,000+36,450

=$168,300

C. Variable Cost per Unit sold = direct material cost + direct labor cost + variable manufacturing overhead + sales commission + variable administrative overhead

= 7.05+3.65+1.6+1.5+0.55

=$14.35

D.Total amount of variable cost for 10,000 units = 14.35*10,000

=$143,500

E. Total amount of manufacturing overhead incurred to produce 10,000 units = variable manufacturing overhead + fixed manufacturing overhead

= 1.6*10,000 + 113,400

= $129,400

F. Contribution Margin per Unit sold = selling price per unit - variable cost per Unit

= 21.6 - 14.35

=$7.25 per Unit

G.Direct manufacturing cost when 8,000 units are produced = (7.05+3.65+1.6)*8,000

=$98,400

H.Indirect manufacturing cost incurred = fixed manufacturing overheads = $113,400

I.i incremental manufacting cost will be variable cost per Unit =$14.35